# Mind Your Theory: Theory of Mind Goes Deeper Than Reasoning

### v2

@@ Abstract
Theory of Mind (ToM) capabilities in LLMs have recently become a central object of investigation, sparking debates and discussions. In this position paper, we explore many lines of work in different communities in AI and cognitive science. Inspired by cognitive work, we view ToM tasks as a two-step process: (I) first, determining whether and how to invoke ToM, which includes setting the appropriate Depth of Mentalizing (DoM); and (II) second, applying correct inference given the appropriate DoM. We identify that many works about ToM in LLMs, such as benchmarks and add-on modules, tend to unjustly overlook the first step and focus exclusively on the second one, which can be framed as static logic problems. We support our empirical evidence about the difficulty of the different steps in existing benchmarks. We conclude with suggestions for improved evaluation of ToM capabilities, inspired by dynamic environments used in cognitive tasks in biological agents.
@@
Eitan Wagner1  Nitay Alon1,2  Joseph M. Barnby3  Omri Abend1 1 Hebrew University of Jerusalem  2 MPI for Biological Cybernetics 3 Royal Holloway University of London {eitan.wagner, nitay.alon}@mail.huji.ac.il
@@ 1 Introduction
The ability of Large Language Models (LLMs) to efficiently integrate social information is essential for ensuring AI trust and safety, as well as reasoning about situations that involve multiple agents. This capability relies on inferring and representing other synthetic and biological agents’ beliefs, desires, and intentions, also known as mentalizing, or Theory of Mind (ToM) Premack and Woodruff (1978).
LLMs have been recently regarded as general-purpose reasoning models Brown et al. (2020), and their ToM capabilities have come under scrutiny, particularly as to whether social cognition can emerge purely from associative principles Sap et al. (2022). Notably, Kosinski (2023) claimed that ToM capabilities emerge in post-GPT3 models, like ToM in children Astington and Jenkins (1995). This claim sparked debates on the correct interpretation of these results Ullman (2023); Pi et al. (2024).
@@ 2.2 ToM Benchmarking
Variations in the SA task have been used as a gold standard benchmark for ToM in LLMs Grant et al. (2017); Nematzadeh et al. (2018); Gandhi et al. (2023). Benchmarks like ToMi Le et al. (2019) and OpenToM Xu et al. (2024), improve previous benchmarks with less predictable generation protocols and more natural settings. Rather than vignette-based tasks, some benchmarks use a language-based conversational setting to assess ToM reasoning, such as Mindcraft Bara et al. (2021), FANToM Kim et al. (2023), and NegotiationToM Chan et al. (2024). Some benchmarks extend the SA task to more complex objectives. This includes higher-order ToM, as in HiToM Wu et al. (2023); additional perception inference, as in Percept-ToMi and Percept-FANToM Jung et al. (2024); and epistemic logic conclusions Sileo and Lernould (2023). Additional extensions derive questions from common ground annotations, as in CommonToM (Soubki et al., 2024) or from other modalities, as in MMToM-QA (Jin et al., 2024). lity of these tasks in assessing ToM in LLMs is still a matter of debate. GPT3 showed difficulties engaging with ToMi Sap et al. (2022), whereas more complex models have been argued to show more success Kosinski (2023); Gandhi et al. (2023). Nevertheless, slight alterations to task structure can alter the performance Ullman (2023); Shapira et al. (2024); Ma et al. (2023a); Chen et al. (2024); Nickel et al. (2024). The theoretical implications of these effects are therefore still under debate Pi et al. (2024). Other criticisms of current benchmarks are the lack of transparency in the training set used to compare LLMs with human-like performance Ma et al. (2023b); Shapira et al. (2024).
In particular, all these benchmarks are static in the sense that they deal with pure observation rather than the evolution of the moment-to-moment model during interaction Ma et al. (2023b). They therefore only distantly resemble the application of ToM capabilities in real-world settings. Additionally, these benchmarks assume that there is always a single correct answer, rather than a range of possible beliefs and behaviors.
@@ 2.3 ToM Add-ons
There are various types of suggested add-ons. Appropriate prompting techniques have shown improved scores on classic ToM tasks Tan et al. (2024); Jung et al. (2024); Huang et al. (2024); Wilf et al. (2024). Explicit symbolic modules can also be updated based on input and improve accuracy Qiu et al. (2024); Hou et al. (2024). Others add post-model decoding methods Sclar et al. (2023) or symbolic planners Jin et al. (2024) to improve performance.
Figure 2: Hierarchical Theory of Mind (ToM). Depth of Mentalizing (DoM) indicates how much an agent (blue) recursively nests beliefs about their partner (orange). At DoM(-1) the agent only considers their own policy ($\pi$). DoM(0) beliefs consider $\pi$ and a partner’s $\pi$. DoM(1) beliefs consider DoM(0) and the beliefs of the partner about the agent’s $\pi$.
@@ 3.1 Invoking DoM
Even with properly invoked ToM reasoning, an agent may still suffer from maladaptive DoM. For example, remaining fixed in a shallow DoM may miss malicious intent from other agents Sarkadi et al. (2019). On the other hand, invoking hierarchical DoM during a cooperative problem can induce false beliefs of harm Alon et al. (2024). Moreover, as described above, when interacting with agents with varying DoM levels (Camerer et al., 2003), a “correct” DoM level may be ill-defined.
@@ 3.2 Revisiting Existing Work
With current vignette-based reasoning tasks, the necessity of ToM is made obvious. Thus, current benchmarks practically address step II of our formulation, thus precluding the need to use ToM or even reason about self vs. other distinctions. Even as challenging benchmarks are introduced, the focus is still on non-interactive presentations with templatized questions, in which the exact required DoM level, $k$, is rather obvious from the question (see Appendix A for a case study on two challenging benchmarks). The problems these benchmarks address are therefore logical ones, which, while challenging, ignore crucial aspects of ToM.
@@ 4 Discussion and Conclusion
In this work, we cast ToM in LLMs as an interdisciplinary topic. We argue that problems that existing NLP benchmarks and AI works typically address are non-interactive logical problems. While many of the benchmarks are valuable as reasoning tests and may provide insights into the performance of LLMs, they do not address the core issue of ToM in LLMs: interaction, and social adaptation.
We argue that a better understanding of ToM in LLMs requires: (1) benchmarks that test whether DoM is correctly invoked or not, (2) evaluation for ToM in interactive settings where the model is an active agent, and (3) evaluation for the appropriate degree of DoM, given a social context (cooperative vs competitive).
Some works introduce ToM tasks for LLMs or formal models in interactive domains Sclar et al. (2022); Li et al. (2023); Ying et al. (2025), albeit with simplified contexts. We believe that a meaningful understanding of ToM capabilities should be through works like these with additional domains and contexts. These tests should cover a variety of ToM complexities (e.g., with decisions about the appropriate DoM), social contexts (e.g., with varying levels of cooperation), and ToM behavior of other agents (e.g., heterogeneous DoM levels in the population).
@@
We also note that the ToM literature in cognitive science is vast and often conflicting. The components of ToM are actively debated and thus our recommendations must be interpreted in light of this.
@@ Acknowledgments
This work was supported in part by the Israel Science Foundation (grant no. 2424/21)

### v3

@@ Abstract
Theory of Mind (ToM) capabilities in LLMs have recently become a central object of investigation, sparking debates and discussions. In this position paper, we explore many lines of work in different communities in AI and cognitive science. Inspired by cognitive work, we view ToM tasks as a two-step process: (I) first, determining whether and how to invoke ToM, which includes setting the appropriate Depth of Mentalizing (DoM); and (II) second, applying correct inference given the appropriate DoM. We identify that many works about ToM in LLMs, such as benchmarks and add-on modules, tend to unjustly overlook the first step and focus exclusively on the second one, which can be framed as a logic-reasoning task. We support our distinction with empirical evidence about the difficulty of the different steps in existing benchmarks. We conclude with suggestions for improved evaluation of ToM capabilities, inspired by dynamic environments used in cognitive tasks in biological agents.
@@
Eitan Wagner1  Nitay Alon1,2  Joseph M. Barnby3,4  Omri Abend1 1 Hebrew University of Jerusalem  2 MPI for Biological Cybernetics 3 Centre for AI and Machine Learning, AU  4 IoPPN, King’s College London {eitan.wagner, nitay.alon}@mail.huji.ac.il
@@ 1 Introduction
The ability of Large Language Models (LLMs) to efficiently integrate social information is essential for ensuring AI trust and safety, as well as for reasoning about situations involving multiple agents. This capability, also known as mentalizing, or Theory of Mind (ToM) Premack and Woodruff (1978), relies on inferring and representing otheragents’ beliefs, desires, and intentions.
LLMs have been recently regarded as general-purpose reasoning models Brown et al. (2020), and their ToM capabilities have come under scrutiny, particularly as to whether social cognition can emerge purely from associative principles Sap et al. (2022). Notably, Kosinski (2023) claimed that ToM capabilities emerge in post-GPT3 models, like ToM in children Astington and Jenkins (1995). This claim sparked debates on the correct interpretation of these results Ullman (2023); Pi et al. (2024); Strachan et al. (2024).
@@ 2.2 ToM Benchmarking
Variations in the SA task have been used as a gold standard benchmark for ToM in LLMs Grant et al. (2017); Nematzadeh et al. (2018); Gandhi et al. (2023). Benchmarks like ToMi Le et al. (2019) and OpenToM Xu et al. (2024), improve previous benchmarks with less predictable generation protocols and more natural settings. Rather than vignette-based tasks, some benchmarks use a language-based conversational setting to assess ToM reasoning, such as Mindcraft Bara et al. (2021), FANToM Kim et al. (2023), and NegotiationToM Chan et al. (2024). Some benchmarks extend the SA task to more complex objectives. This includes higher-order ToM, as in HiToM Wu et al. (2023); additional perception inference, as in Percept-ToMi and Percept-FANToM Jung et al. (2024); and epistemic logic conclusions Sileo and Lernould (2023). Additional extensions derive questions from common ground annotations, as in CommonToM (Soubki et al., 2024) or from other modalities, as in MMToM-QA (Jin et al., 2024).
The utility of these tasks in assessing ToM in LLMs is still a matter of debate. GPT3 showed difficulties engaging with ToMi Sap et al. (2022), whereas more complex models have been argued to show more success Kosinski (2023); Gandhi et al. (2023). Nevertheless, slight alterations to task structure or details can alter the performance Ullman (2023); Shapira et al. (2024); Ma et al. (2023a); Chen et al. (2024); Nickel et al. (2024). The theoretical implications of these effects are therefore still under debate Pi et al. (2024). Other criticisms of current benchmarks are due to the lack of transparency of training sets, hindering the ability to compare LLMs with human performance Ma et al. (2023b); Shapira et al. (2024).
In particular, all these benchmarks are static in the sense that they deal with pure observation rather than the evolution of the moment-to-moment model during interaction Ma et al. (2023b). They therefore only distantly resemble the application of ToM capabilities in real-world settings. Additionally, these benchmarks assume that there is always a single correct answer, for which all the relevant information is given, rather than a range of possible beliefs and behaviors.
@@ 2.3 ToM Add-ons
There are various types of suggested add-ons. Task-specific prompting techniques have shown improved scores on classic ToM tasks Tan et al. (2024); Jung et al. (2024); Huang et al. (2024); Wilf et al. (2024). Explicit symbolic modules can also be updated based on input and improve accuracy Qiu et al. (2024); Hou et al. (2024). Others add post-model decoding methods Sclar et al. (2023) or symbolic planners Jin et al. (2024) to improve performance.
Figure 2: Hierarchical Dynamic Theory of Mind (ToM). Depth of Mentalizing (DoM) indicates how much an agent (blue) recursively nests beliefs about their partner (orange). At DoM(-1) the agent only considers their own policy ($\pi$). DoM(0) beliefs consider $\pi$ and a partner’s $\pi$. DoM(1) beliefs consider DoM(0) and the beliefs of the partner about the agent’s $\pi$.
@@ 3 What Can We Learn From Biological Agents?
More generally, the first step corresponds to the act of “theorizing” about the agents’ belief structure, and the second step corresponds to the application of this theory. The theory can be reassessed when new information is available, forming a dynamic theory.
@@ 3.1 Invoking DoM
Even with properly invoked ToM reasoning, an agent may still suffer from maladaptive DoM. For example, remaining fixed in a shallow DoM may miss malicious intent from other agents Sarkadi et al. (2019). On the other hand, invoking hierarchical DoM during a cooperative problem can induce false beliefs of harm Alon et al. (2024). Moreover, as described above, when interacting with agents with varying DoM levels (Camerer et al., 2003), a “correct” DoM level may be ill-defined. Arguably, modeling DoM involves estimating a distribution, as opposed to question-answering scenarios which involve predicting a single response (Wagner and Abend, 2025).
@@ 3.2 Revisiting Existing Work
With current vignette-based reasoning tasks, the necessity of ToM is made obvious. Thus, current benchmarks practically address step II of our formulation, thus precluding the need to use ToM or even reason about self vs. other distinctions. Even as challenging benchmarks are introduced, the focus is still on non-interactive presentations with templatized questions, in which the exact required DoM level, $k$, is rather obvious from the question. The problems these benchmarks address are therefore logical ones, which, while challenging, ignore crucial aspects of ToM.
We experimentally support this claim on two challenging false-belief benchmarks, HiToM Wu et al. (2023) and FANToM Kim et al. (2023). In addition to the belief-related questions arising from the described scenarios, we add a task of predicting the required ToM order for each question. The ToM order is recorded as metadata as it guides the generation process of the scenarios, providing gold-truth labels. Arguably, accurately identifying the required ToM order while incorrectly answering questions about the beliefs, shows that the main challenge is not in the ToM invocation step. Our results show this is indeed the case: while GPT-3.5-turbo struggles even with this task, GPT-4o-mini performs well, and GPT-4o is nearly perfect. Chain-of-thought (CoT) prompting gives some improvement. See Appendix A for full details of the experiments and results.
@@
Why is this important?
The influence of cognitive science is evidenced in many ToM benchmark papers. Many works (e.g., Kim et al., 2023; Jin et al., 2024) mention and adopt psychology principles that are required when designing a task to validate ToM. Additionally, the active debate about emerging ToM capabilities (§2.2) is highly connected to the assumption that ToM capabilities reflect more than a specific logical task. In general, there seems to be an underlying assumption that ToM capabilities in models should cover the same tasks that humans perform by using ToM. Many works (e.g., Ullman, 2023; Shapira et al., 2024) question the generalization and robustness of LLM performance within a specific ToM task. Our formalization further suggests that a crucial cognitive part of ToM behavior in humans, which has its unique properties, is rarely addressed in common LLM literature.
@@ 4 Discussion and Conclusion
In this work, we cast ToM in LLMs as an interdisciplinary topic. We argue that the problems addressed by existing NLP benchmarks and AI works are essentially non-interactive logical problems. While many of the benchmarks are valuable as reasoning tests and may provide insights into the performance of LLMs, they do not address the core issue of ToM in LLMs: interaction, and social adaptation.
We argue that a better understanding of ToM in LLMs requires: (1) benchmarks that test whether DoM is correctly invoked or not, (2) evaluation for ToM in interactive settings where the model is an active agent, and (3) distributional assessment of DoM dynamics, given a social context (cooperative vs competitive).
The requirement for social dynamism is expressed in cognitive science and computational psychology. Neural outcomes show distinct representations when considering non-social versus social outcomes Zhang and Gläscher (2020), demanding a theoretical distinction that captures social dynamics sufficiently Cushman (2024); Jara-Ettinger (2019); Barnby et al. (2023, 2024). We believe this distinction should also be reflected in the way we benchmark models for ToM.
Wang et al. (2025), argue in favor of dynamic, user-centric benchmarks for ToM evaluation. Our work extends this call, emphasizing the role of dynamic settings for a phase in the ToM reasoning that is commonly overlooked. Indeed, some previous works introduce ToM tasks for LLMs or formal models in interactive domains Sclar et al. (2022); Li et al. (2023); Ying et al. (2025). However, these benchmarks are limited in their scope as they tend to model simplified fixed worlds. We believe that a meaningful understanding of ToM capabilities should be achieved by extending works like these with additional domains and contexts. Whereas a single scenario may not fully capture the range of behavior, constructing multiple diverse ones should lead to better coverage of ToM complexities. This should include situations without full context (e.g., without explicit information about the requested), social situations (e.g., with varying levels of cooperation), and ToM behavior of other agents (e.g., heterogeneous DoM levels in the population).
@@
We also note that the ToM literature in cognitive science is vast and often conflicting. The components of ToM are actively debated; thus our recommendations must be interpreted in light of this.
@@ Acknowledgments
This work was supported in part by the Israel Science Foundation (grant no. 2424/21). NA is supported by the Max Planck Society and the Humboldt Foundation.